# revtex4-2Repair the float

### v1

@@ Abstract
Machine learning offers a largely unexplored avenue for improving noisy disordered devices in physics using automated algorithms. Through simulations that include disorder in physical devices, particularly quantum devices, there is potential to learn about disordered landscapes and subsequently tune devices based on those insights. In this work, we introduce a novel methodology that employs machine learning, specifically convolutional neural networks (CNNs), to discern the disorder landscape in the parameters of the disordered extended Hubbard model underlying the semiconductor quantum dot spin qubit architectures. This technique takes advantage of experimentally obtainable charge stability diagrams from neighboring quantum dot pairs, enabling the CNN to accurately identify disorder in each parameter of the extended Hubbard model. Remarkably, our CNN can process site-specific disorder in Hubbard parameters, including variations in hopping constants, on-site potentials (gate voltages), and both intra-site and inter-site Coulomb terms. This advancement facilitates the prediction of spatially dependent disorder across all parameters simultaneously with high accuracy ($R^{2}>0.994$) and fewer parameter constraints, marking a significant improvement over previous methods that were focused only on analyzing on-site potentials at low coupling. Furthermore, our approach allows for the tuning of five or more quantum dots at a time, effectively addressing the often-overlooked issue of crosstalk. Not only does our method streamline the tuning process, potentially enabling fully automated adjustments, but it also introduces a "no trust" verification method to rigorously validate the neural network’s predictions. Ultimately, this work aims to lay the groundwork for generalizing our method to tackle a broad spectrum of physical problems. In particular, our work establishes that the microscopic parameters controlling the semiconductor quantum dot quantum computing platforms can be uniquely determined in an automated manner by using a CNN based machine learning technique using only the measured charge stability diagrams as the input.
@@
Two current areas of high scientific and technological (in fact, even public) interest are neural networks and quantum computation. While neural networks (or more generally, machine learning) have progressed much more rapidly recently, quantum computation has also been making significant progress. Machine learning provides a tool that could potentially speed up advances in the development of quantum information hardware. The issue of interest to us is the inevitable presence of unknown and unintentional random disorder in the quantum devices, which necessitates repeated finetuning and re-calibration of the devices in order to obtain meaningful and reproducible experimental results. We propose that neural network based machine learning could provide considerable help in this tuning problem in an unknown disorder landscape. In particular, for a large class of physical problems, there exists a sufficiently accurate physical model that is disordered in an unknown but potentially predictable manner. For this class of problems, it may be possible through theoretical simulations to generate enough training data that allows a neural network to determine the disorder landscape of the problem and then feasibly correct it.
@@
The semiconductor microelectronics industry is mature, because of its key role in all modern technologies, thereby enabling the integration of quantum dot technology into the pre-existing infrastructure. Additionally, quantum dots can be created sufficiently small in size consistent with the existing microelectronics industry, allowing for scalability to large number of qubits required to perform advanced quantum algorithms [24]. This is a huge potential scalability and manufacturing advantage driving quantum computing research in the quantum dot based qubit platforms. In fact, recently Si quantum dot qubits have been fabricated in advanced manufacturing fabs where industrial semiconductor wafers are made, establishing the microelectronic manufacturing compatibility and potential of quantum dot architectures [16, 35]. Semiconductor quantum dots in their quantum computing architecture, where the two-qubit gates are operated by inter-dot exchange coupling, are naturally described by the Hubbard model in the low-temperature, strong-interaction regime, which facilitates their numerical simulation [25, 13, 10, 5, 20]. This is typically set up as a Hubbard model in the form of a chain (or plaquette) of sites (with each quantum dot being a site) of some length. Disorder significantly impedes the use of quantum dots for computational purposes, making the control of disorder one of the most important challenges to overcome in utilizing quantum dots as a platform for quantum computation. The disorder within quantum dots often manifests as errors in each Hubbard model parameter, which are generally unknown, vary between sites, and must be corrected before utilizing the dots for computation. Automating the control of such disordered quantum dot structures in order to carry out precise gate operations is a serious practical problem since manual control can only work for a system of a few dots [36, 37]. Scaling up to a many-dot quantum computing platform requires a decisive resolution of the tuning problem in the form of an automated procedure.
@@
In previous work [10], a method was developed to determine site-dependent disorder in the gate voltage for three quantum dots, using a essentially a by-hand trial and error algorithm to correct the disorder in the gate voltages of the dot chain. This methodology has been invaluable for experimentalists aiming to employ quantum dots for quantum computational purposes. Specifically, this approach utilizes charge stability diagrams, which are easily produced and serve as the primary means for understanding the properties of these devices. For quantum dots to be viable for quantum computation, it is crucial that their Hubbard parameters are tuned to specific values so that the Hamiltonian is precisely defined (which is an essential condition for error-free initialization and gate operations). However, this becomes a challenging task as these parameters are prone to natural drift over time due to environmental factors, noise and disorder, necessitating periodic re-tuning through human intervention. The laborious process of re-tuning quantum dots is both unwelcome and time-consuming, typically requiring the examination of numerous charge stability diagrams by-hand to identify the optimal corrections for parameter drift. Current methodologies [7] also limit adjustments to only two quantum dots at a time, since attempting to calibrate all the Hubbard model parameters simultaneously for a sequence of quantum dots is impractical for an individual. Clearly, a much better automatic control must be developed for scaling up the quantum dot structures since calibration and re-tuning individually is out of question for more than 2-3 dots.
@@
Machine learning presents an opportunity for improvement. Neural networks, through the use of large amounts of data, have been able to make connections and thus predictions rapidly for tasks that would require impossibly extensive human effort (e.g. image processing). In physics, machine learning techniques have been utilized to provide valuable insights into disordered problems, including potentially determining the entire disorder landscape [28]. When a sufficiently accurate theoretical model exists (i.e. a precise Hamiltonian such as the Hubbard model for quantum dot qubits)- even if it cannot be directly applied due to disorder within its parameters, which take an understandable (i.e. not arbitrary) but unknown (and random) form in the Hamiltonian—it is possible to generate large amounts of training data from simulations. This training data can often enable a deep neural network to find the disordered parameters in a way that is applicable to experimental devices. Furthermore, if the device can be modeled accurately, one can conduct verification tests beyond the neural network’s input to validate its predictions.
@@
Beyond this, we also introduce a "no trust" method for verifying the output of our neural network, which offers a significant advantage over other neural network approaches by preventing the possibility of hallucination and enabling conclusive experimental empirical confirmation of the model’s predictions. Finally, we conclude by generalizing our approach, discussing how our method can be applied to solve disorder in a broad range of physics problems while utilizing the "no trust" confirmation scheme we have implemented in this work and previously. The paper is organized as follows:
@@ Methods
      Current Tuning Process: Briefly describes the process for tuning quantum dots using charge stability diagrams.
@@ Results
      No Trust Verification: Details a method to verify machine learning results without relying on the neural network output, addressing the issue of hallucination.
@@ I.1 Model
The objective is to ascertain the deviations in each Hubbard model parameter, which effectively determines the parameters of the disordered Hamiltonian. The system is considered at a finite (low) temperature $\frac{k_{B}}{\beta}=0.005\langle U\rangle$ (with $\langle U\rangle$ being the average U value over the sites), with site-dependent chemical potentials $\mu_{i}$ being variable. Our measurements are based on expectation values $\langle n_{i}\rangle=\text{Tr}(\rho n_{i})$, where $\rho=e^{-\beta(H-\mu_{i}n_{i})}$, and are taken across an array of different $\mu_{i}$ values. Experimentally, these measurements are facilitated by applying a site-dependent gate voltage. All sites are measured simultaneously, resulting in an occupation vector $\vec{n}=(\langle n_{1}\rangle,\cdots,\langle n_{N}\rangle)$. Numerically, the simulation was carried out through exact diagonalization using the Quspin package [31]. To account for thermal effects, the bottom 10 lowest energy states were retained for generating $\rho$, an amount that is more than adequate (for describing the typical experimental situation) given the low experimental temperature setting. Note that as long as the temperature is low, its precise value is irrelevant.
@@ I.2 Current Tuning Process
The current method for determining the Hubbard model parameters for quantum dots involves assessing the parameters for each pair of dots by examining their charge stability diagrams. This approach requires assuming that the effects of other dots can be either removed or rendered negligible. Charge stability diagrams depict the stable electronic configurations (or expected occupation numbers) across different sites as a function of the chemical potentials (or experimentally, gate voltages). In the scenario of two quantum dots, the Hubbard parameters’ influence on the stability diagrams behaves in a predictable manner, making it possible to deduce these parameters. For example, the parameter t visually curves the boundaries within the charge stability diagram. The value of t between two quantum dots can be inferred from the boundary between the (0,0) occupation phase and the (0,1) or (1,0) occupation phases. The boundary is represented by the equation $\mu_{1}*\mu_{2}=t^{2}$, leading to the triple point labeled A, where $\mu_{1}(A)=\mu_{2}(A)=-t$ [32]. Assuming t is sufficiently small (which is the generic situation), $U_{i}$ can initially be approximated within the stability diagram, based on the energy needed to add a single electron to the i-th dot. If it is known that $U_{1}=U_{2}=U$, then $V_{12}$ for two quantum dots can be determined from the position of a triple point within the charge stability diagram [30]. Triple points are intersections of three distinct phases. The necessary triple point, labeled B, involves [(1,0),(0,1),(1,1)] occupations and occurs at $\mu_{1}(B)=\mu_{2}(B)=t+\frac{(U+V_{12})}{2}-\sqrt{4t^{2}+\frac{(U-V_{12})^{2}}{4}}$. In situations where $U_{1}\neq U_{2}$, or when t is not small enough to simplify the $U_{i}$ determination, numerical fitting should be done manually to determine $V_{12}$, highlighting a primary limitation in current methodologies [5]. For more details on this background, we refer the reader to Refs. [10, 5, 30].
The site-dependent disorder, $\tilde{\epsilon}_{i}$, has been investigated for tuning three quantum dots in a previous work [10], where a method of least squares with phase diagram transition lines (identified using an edge-finding algorithm) is employed to fit the parameters. This method proves effective but is limited to small $t/U<0.15$, as it fails at larger coupling values because the phase transition lines cease to be sufficiently straight. In summary, the current method for determining the Hubbard model parameters relies heavily on manual fitting beyond a certain point, using two or at most three quantum dots, which is notably inefficient.
@@ I.3 Training Data
The training data for the machine learning model consists of performing occupation expectation $\vec{n}$ measurements across various measurement configurations. Each configuration specifies a list of $\mu_{i}$ for every site. However, for practical purposes, within each nearest neighbor pair of quantum dots $i$ and $i+1$, we assign the parameters as follows: $\mu_{i}=\mu_{i}$, $\mu_{i+1}=-\mu_{i}$, and for any $\mu_{j\neq i}=\mu$. We then independently vary $\mu$ and $\mu_{i}$ for all sites in our model and across every pair of nearest neighbor sites $i$ and $i+1$, using periodic boundary conditions. (Using open boundary conditions makes the machine learning procedure somewhat easier numerically, with similar results– the periodic/open boundary conditions physically correspond to 2D/1D qubit configurations.) Specifically, $\mu$ is adjusted between $[-0.5W,2.5W]$ with 50 steps, and $\mu_{i}$ is varied between $[-0.6\langle U\rangle,+0.6\langle U\rangle]$ with 20 steps, where $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ acts as the normalization factor. These normalization factors are based on their deviation-free values, consistent across all disorder realizations, although this adjustment is merely a linear factor and not essential. The various measurement configurations are recorded in a matrix K, where each row records a specific setup of experimentally tunable parameters, labeled as j. The structure of the K matrix is presented below:
@@
For each measurement configuration (or row $K^{j}$), the expected occupation number of each quantum dot is used to form an input matrix $X$. This $X$ matrix, formed from experimental measurements, will be utilized by the neural network to determine the Hubbard parameters. In our case, of course, these ’experimental measurements’ producing the data are generated by the actual simulations of the disordered Hubbard model.
@@
Putting everything together for clarity, our input X is a series of a version of charge stability diagrams. The neural network takes in a series of nearest neighbor charge stability diagrams to determine the Hubbard model parameters for a ring of quantum dots.
@@ I.4 Machine Learning
Two different but very similar architectures were employed in the current work: one with 2D convolutional layers and another with 3D convolutional layers. These setups, depicted in Fig. 2, will be referred to as the 3D type and the 2D type, respectively. In both types, the neural network consists of two sets of convolutional layers, each containing two layers. In the 2D (3D) type, the first set includes two 2D (3D) convolutional layers with 64 filters and a kernel size of 3x3 (3x3x3), while the second set is identical but with 128 filters. A 2D (3D) max pooling layer with a size of 2x2 (2x2x2) is placed between each set. Following the convolutional sets in both types are two dense layers with 1000 nodes each and a final dense layer equal to the output size. This relatively simple neural network configuration is notably efficient in predicting the Hubbard parameters. For the 2D type, the input was in the form of charge stability diagrams, as previously mentioned, but these were reshaped into a 2D grid based on the $\mu$ and $\mu_{i}^{j}$ values, where the channels represent the expected occupation values of different quantum dots. However, this setup does not include the site number $i_{site}$ in the input. To incorporate this parameter, the 2D grid for each different site pair was appended sequentially to the 2D grid, resulting in input data that took the form of a grid of occupation numbers based on the $\mu$ and $\mu_{i}^{j}$ values, where $i=1,2\cdots$, all appended together along the $\mu^{j}$ axis. This peculiar setup was chosen for the 2D type to facilitate the use of a 2D CNN and to make the input more symmetrical. In the 3D type, this appending was unnecessary as the convolution spanned all three axes: $\mu$, $\mu_{i}^{j}$, and $i_{site}$. The 2D type architecture was utilized for predictions involving deviations in a single parameter, whereas the 3D type architecture was applied to predict deviations in multiple parameters. For both types, 90% of the generated disorder realizations were used for training, and the remaining 10% were reserved for testing and validation. The validity of our model was assessed by calculating the $R^{2}$ (coefficient of determination) and the root mean squared (RMS) error for our predictions on the validation data. The training (90%) and testing (10%) data were selected randomly with no bias.

### v2

@@ Abstract
In this work, we introduce a novel methodology that employs machine learning, specifically convolutional neural networks (CNNs), to discern the disorder landscape in the parameters of the disordered extended Hubbard model underlying the semiconductor quantum dot spin qubit architectures. Our approach, which we demonstrate with numerically simulated devices takes advantage of currently used experimentally obtainable charge stability diagrams from neighboring quantum dot pairs, enabling the CNN to accurately identify disorder in each parameter of the extended Hubbard model. Remarkably, our CNN can process site-specific disorder in Hubbard parameters, including variations in hopping constants, on-site potentials (gate voltages), and both intra-site and inter-site Coulomb terms. This advancement facilitates the prediction of spatially dependent disorder across all parameters simultaneously with high accuracy ($R^{2}>0.994$) and fewer parameter constraints, marking a significant improvement over previous methods that were focused only on analyzing on-site potentials at low coupling. Furthermore, our approach allows for the tuning of five or more quantum dots at a time, effectively addressing the often-overlooked issue of crosstalk. Not only does our method streamline the tuning process, potentially enabling fully automated adjustments, but it also introduces a "no trust" verification method to rigorously validate the neural network’s predictions. Ultimately, this work aims to lay the groundwork for generalizing our method to tackle a broad spectrum of physical problems.
@@
Two current areas of high scientific and technological (in fact, even public) interest are neural networks and quantum computation. While neural networks have progressed much more rapidly recently, quantum computation has also been making significant progress. Machine learning provides a tool that could potentially speed up advances in the development of quantum information hardware. The issue of interest to us is the inevitable presence of unknown and unintentional random disorder in the quantum devices, which necessitates repeated finetuning and re-calibration of the devices in order to obtain meaningful and reproducible experimental results. We propose that neural network based machine learning could provide considerable help in this tuning problem in an unknown disorder landscape. In particular, for a large class of physical problems, there exists a sufficiently accurate physical model that is disordered in an unknown but potentially predictable manner. For this class of problems, it may be possible through theoretical simulations to generate enough training data that allows a neural network to determine the disorder landscape of the problem and then feasibly correct it.
@@
The semiconductor microelectronics industry is mature, because of its key role in all modern technologies, thereby enabling the integration of quantum dot technology into the pre-existing infrastructure. Additionally, quantum dots can be created sufficiently small in size consistent with the existing microelectronics industry, allowing for scalability to large number of qubits required to perform advanced quantum algorithms [18]. This is a huge potential scalability and manufacturing advantage driving quantum computing research in the quantum dot based qubit platforms. In fact, recently Si quantum dot qubits have been fabricated in advanced manufacturing fabs where industrial semiconductor wafers are made, establishing the microelectronic manufacturing compatibility and potential of quantum dot architectures [19, 20]. Semiconductor quantum dots in their quantum computing architecture, where the two-qubit gates are operated by inter-dot exchange coupling, are naturally described by the Hubbard model in the low-temperature, strong-interaction regime, which facilitates their numerical simulation [21, 22, 14, 23, 24]. This is typically set up as a Hubbard model in the form of a chain (or plaquette) of sites of some length. Disorder significantly impedes the use of quantum dots for computational purposes, making the control of disorder one of the most important challenges to overcome in utilizing quantum dots as a platform for quantum computation. The disorder within quantum dots often manifests as errors in each Hubbard model parameter, which are generally unknown, vary between sites, and must be corrected before utilizing the dots for computation. Automating the control of such disordered quantum dot structures in order to carry out precise gate operations is a serious practical problem since manual control can only work for a system of a few dots [25, 26]. Scaling up to a many-dot quantum computing platform requires a decisive resolution of the tuning problem in the form of an automated procedure.
@@
In previous work [14], a method was developed to determine site-dependent disorder in the gate voltage for three quantum dots, using a essentially a by-hand trial and error algorithm to correct the disorder in the gate voltages of the dot chain. This methodology has been invaluable for experimentalists aiming to employ quantum dots for quantum computational purposes. Specifically, this approach utilizes charge stability diagrams, which are easily produced and serve as the primary means for understanding the properties of these devices. There have also been other works making use of machine learning for this platform [27, 28, 29]. We mention that the main point of our work, a theoretical machine learning based automation of controlling charge stability in semiconductor quantum dot qubits, has recently been identified as one of the central problems preventing progress in semiconductor based solid state qubits [25, 26]. For quantum dots to be viable for quantum computation, it is crucial that their Hubbard parameters are tuned to specific values so that the Hamiltonian is precisely defined (which is an essential condition for error-free initialization and gate operations). However, this becomes a challenging task as these parameters are prone to natural drift over time due to environmental factors, noise and disorder, necessitating periodic re-tuning through human intervention. The laborious process of re-tuning quantum dots is both unwelcome and time-consuming, typically requiring the examination of numerous charge stability diagrams by-hand to identify the optimal corrections for parameter drift. Current methodologies [30] also limit adjustments to only two quantum dots at a time, since attempting to calibrate all the Hubbard model parameters simultaneously for a sequence of quantum dots is impractical for an individual. Clearly, a much better automatic control must be developed for scaling up the quantum dot structures since calibration and re-tuning individually is out of question for more than 2-3 dots.
@@
Machine learning presents an opportunity for improvement. Neural networks, through the use of large amounts of data, have been able to make connections and thus predictions rapidly for tasks that would require impossibly extensive human effort (e.g. image processing). In physics, machine learning techniques have been utilized to provide valuable insights into disordered problems, including potentially determining the entire disorder landscape [31]. When a sufficiently accurate theoretical model exists - even if it cannot be directly applied due to disorder within its parameters, which take an understandable but unknown form in the Hamiltonian—it is possible to generate large amounts of training data from simulations. This training data can often enable a deep neural network to find the disordered parameters in a way that is applicable to experimental devices. Furthermore, if the device can be modeled accurately, one can conduct verification tests beyond the neural network’s input to validate its predictions.
@@
Beyond this, we also introduce a "no trust" method for verifying the output of our neural network, which offers a significant advantage over other neural network approaches by preventing the possibility of hallucination and enabling conclusive experimental empirical confirmation of the model’s predictions. Finally, we conclude by generalizing our approach, discussing how our method can be applied to solve disorder in a broad range of physics problems while utilizing the "no trust" confirmation scheme we have proposed in this work and previously. The paper is organized as follows:
@@ Methods
      Current Tuning Process: Provides background briefly describing the currently used process for tuning quantum dots using charge stability diagrams.
@@ Results
      No Trust Verification Scheme: Details a method to verify machine learning results without relying on the neural network output, addressing the issue of hallucination.
@@ I.1 Model
The objective is to ascertain the deviations in each Hubbard model parameter, which effectively determines the parameters of the disordered Hamiltonian. The system is considered at a finite (low) temperature $\frac{k_{B}}{\beta}=0.005\langle U\rangle$ (with $\langle U\rangle$ being the average U value over the sites), with site-dependent chemical potentials $\mu_{i}$ being variable. Our measurements are based on expectation values $\langle n_{i}\rangle=\text{Tr}(\rho n_{i})$, where $\rho=e^{-\beta(H-{\sum_{i}}\mu_{i}n_{i})}$, and are taken across an array of different $\mu_{i}$ values. Experimentally, these measurements are facilitated by applying a site-dependent gate voltage. All sites are measured simultaneously, resulting in an occupation vector $\vec{n}=(\langle n_{1}\rangle,\cdots,\langle n_{N}\rangle)$. Numerically, the simulation was carried out through exact diagonalization using the Quspin package [33]. To account for thermal effects, the bottom 10 lowest energy states were retained for generating $\rho$, an amount that is more than adequate (for describing the typical experimental situation) given the low experimental temperature setting. Note that as long as the temperature is low, its precise value is irrelevant. Even in the case of larger systems with small coupling constant, a lower T say $\frac{k_{B}}{\beta}=0.0026\langle U\rangle$, which is experimentally achievable [14], should be sufficient to maintain an effectively pure state for significantly larger system sizes than we consider.)
@@ I.2 Current Tuning Process
The currently used method for determining the Hubbard model parameters for quantum dots involves assessing the parameters for each pair of dots by examining their charge stability diagrams. This approach, unlike ours, requires assuming that the effects of other dots can be either removed or rendered negligible. Charge stability diagrams depict the stable electronic configurations (or expected occupation numbers) across different sites as a function of the chemical potentials (or experimentally, gate voltages). In the scenario of two quantum dots, the Hubbard parameters’ influence on the stability diagrams behaves in a predictable manner, making it possible to deduce these parameters. For example, the parameter t visually curves the boundaries within the charge stability diagram. The value of t between two quantum dots can be inferred from the boundary between the (0,0) occupation phase and the (0,1) or (1,0) occupation phases. Experimentally, these boundaries are peaks in the charge addition spectrum. The boundary is represented by the equation $\mu_{1}*\mu_{2}=t^{2}$, leading to the triple point labeled A, where $\mu_{1}(A)=\mu_{2}(A)=-t$ [34]. Assuming t is sufficiently small (which is the generic situation), $U_{i}$ can initially be approximated within the stability diagram, based on the energy needed to add a single electron to the i-th dot. If it is known that $U_{1}=U_{2}=U$, then $V_{12}$ for two quantum dots can be determined from the position of a triple point within the charge stability diagram [35]. Triple points are intersections of three distinct phases. The necessary triple point, labeled B, involves [(1,0),(0,1),(1,1)] occupations and occurs at $\mu_{1}(B)=\mu_{2}(B)=t+\frac{(U+V_{12})}{2}-\sqrt{4t^{2}+\frac{(U-V_{12})^{2}}{4}}$. ( See Fig. 1) In situations where $U_{1}\neq U_{2}$, or when t is not small enough to simplify the $U_{i}$ determination, numerical fitting should be done manually to determine $V_{12}$, highlighting a primary limitation in current methodologies [23]. For more details on this background, we refer the reader to Refs. [14, 23, 35].
The site-dependent disorder, $\tilde{\epsilon}_{i}$, has been investigated for tuning three quantum dots in a previous work [14], where a method of least squares with phase diagram transition lines (identified using an edge-finding algorithm) is employed to fit the parameters. This method proves effective but is limited to small $t/U<0.15$, as it fails at larger coupling values because the phase transition lines cease to be sufficiently straight and $n_{i}$ cease to be good quantum numbers making the boundaries diffuse. In summary, the current method for determining the Hubbard model parameters relies heavily on manual fitting beyond a certain point, using two or at most three quantum dots, which is notably inefficient.
Figure 1: Standard charge stability diagram showing the most probable states calculated at $U_{1}=U_{2}=6$, $t=0.1$, $V_{12}=2$ and $\epsilon_{i}=0$
@@ I.3 Training Data
The training data for the machine learning model consists of performing occupation expectation $\vec{n}$ measurements across various measurement configurations. Each configuration specifies a list of $\mu_{i}$ for every site. However, for practical purposes, within each nearest neighbor pair of quantum dots $i$ and $i+1$, we assign the parameters as follows: $\mu_{i}$, $\mu_{i+1}=-\mu_{i}$, and for any $\mu_{j\neq i}=\mu$. We then independently vary $\mu$ and $\mu_{i}$ for all sites in our model and across every pair of nearest neighbor sites $i$ and $i+1$, using periodic boundary conditions. (Using open boundary conditions makes the machine learning procedure somewhat easier numerically, with similar results– the periodic/open boundary conditions physically correspond to 2D/1D qubit configurations, namely 4 dots with periodic boundaries is equivalent to 2x2 dots with an open boundary.) Specifically, $\mu$ is adjusted between $[-0.5W,2.5W]$ with 50 steps, and $\mu_{i}$ is varied between $[-0.6\langle U\rangle,+0.6\langle U\rangle]$ with 20 steps, where $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ acts as the normalization factor. These normalization factors are based on their deviation-free values, consistent across all disorder realizations, although this adjustment is merely a linear factor and not essential. It should be noted that this measurement space is significantly less than the available observation space, this being sufficient is surprising and determined in this work posteriori. The various measurement configurations are recorded in a matrix K, where each row records a specific setup of experimentally tunable parameters, labeled as j. The structure of the K matrix is presented below:
@@
Where $i^{j}_{site}$ records the site of $\mu_{i}^{j}$. For each measurement configuration (or row $K^{j}$), the expected occupation number of each quantum dot is used to form an input matrix $X$. This $X$ matrix, formed from experimental measurements, will be utilized by the neural network to determine the Hubbard parameters, though in our case, of course, these ’experimental measurements’ producing the data are generated by the actual simulations of the disordered Hubbard model:
@@
Putting everything together for clarity, our input X is a series of a version of charge stability diagrams. The neural network takes in a series of nearest neighbor charge stability diagrams to determine the Hubbard model parameters for a ring of quantum dots. This input is essentially the same as what is already being used to experimentally tune dots without any additional processing [14]. A pristine example of the diagram can be seen in Fig. 2.
@@ I.4 Machine Learning
Two different but very similar architectures are employed in the current work: one with 2D convolutional layers and another with 3D convolutional layers. These setups, depicted in Fig. 3, will be referred to as the 3D type and the 2D type, respectively. In both types, the neural network consists of two sets of convolutional layers, each containing two layers. In the 2D (3D) type, the first set includes two 2D (3D) convolutional layers with 64 filters and a kernel size of 3x3 (3x3x3), while the second set is identical but with 128 filters. A 2D (3D) max pooling layer with a size of 2x2 (2x2x2) is placed between each set. Following the convolutional sets in both types are two dense layers with 1000 nodes each and a final dense layer equal to the output size. This relatively simple neural network configuration is notably efficient in predicting the Hubbard parameters. For the 2D type, the input is in the form of charge stability diagrams, as previously mentioned, but these were reshaped into a 2D grid based on the $\mu$ and $\mu_{i}^{j}$ values, where the channels represent the expected occupation values of different quantum dots. However, this setup does not include the site number $i_{site}$ in the input. To incorporate this parameter, the 2D grid for each different site pair is appended sequentially to the 2D grid, resulting in input data that took the form of a grid of occupation numbers based on the $\mu$ and $\mu_{i}^{j}$ values, where $i=1,2\cdots$, all appended together along the $\mu^{j}$ axis. This peculiar setup was chosen for the 2D type to facilitate the use of a 2D CNN and to make the input more symmetrical. In the 3D type, this appending is unnecessary as the convolution spanned all three axes: $\mu$, $\mu_{i}^{j}$, and $i_{site}$. The 2D type architecture is utilized for predictions involving deviations in a single parameter, whereas the 3D type architecture is applied to predict deviations in multiple parameters. For both types, 90% of the generated disorder realizations were used for training, and the remaining 10% were reserved for testing and validation. The validity of our model is assessed by calculating the $R^{2}$ (coefficient of determination) and the root mean squared (RMS) error for our predictions of the deviations on the validation data. The training (90%) and testing (10%) data were selected randomly with no bias.